OpenStudy (anonymous):

f(x) = 19xe^(1 − x^2) Find the exact location of all the relative and absolute extrema of the function @ganeshie8

ganeshie8 (ganeshie8):

same thing 1) find the first derivative 2) set it equal to 0 and solve x

OpenStudy (anonymous):

19e^(1-x^2) (-2x)?

ganeshie8 (ganeshie8):

you need to use product rule and chainrule

ganeshie8 (ganeshie8):

wolfram gives this http://www.wolframalpha.com/input/?i=%2819xe%5E%281+%E2%88%92+x%5E2%29%29%27

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

so now set it = to 0??

ganeshie8 (ganeshie8):

yes

ganeshie8 (ganeshie8):

\[\large f'(x) = -19e^{1-x^2}(2x^2-1) = 0\]

ganeshie8 (ganeshie8):

obviously 19 is not 0

ganeshie8 (ganeshie8):

can `e^something` ever become 0 ?

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

is it 1/squareroot 2

ganeshie8 (ganeshie8):

you should get +- 1/sqrt(2)

ganeshie8 (ganeshie8):

two points

OpenStudy (anonymous):

so my answer is just 1/sqrt2,-1/sqrt2?

ganeshie8 (ganeshie8):

nope, those are the x values

ganeshie8 (ganeshie8):

plugin each value into f(x) for y coordinate

OpenStudy (anonymous):

for the positive i got 23.4942781075

ganeshie8 (ganeshie8):

doesnt look correct

OpenStudy (anonymous):

>.<

OpenStudy (anonymous):

yea its wrong

OpenStudy (anonymous):

so it is saying that it is wrong

ganeshie8 (ganeshie8):

wolfram is giving y = 22.1506 when x = 1/sqrt(2) right ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

so one critical point is (1/sqrt(2) , 22.1506)

OpenStudy (anonymous):

maximum

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

but it is saying its wrong

ganeshie8 (ganeshie8):

the minimum wold be (-1/sqrt(2) , - 22.1506)

OpenStudy (anonymous):

it still wrong>.<

ganeshie8 (ganeshie8):

try entering `(0.71, 22.15)` for maximum

ganeshie8 (ganeshie8):

online graders are usually dumb you need to spoonfeed them in the exact format they digest

OpenStudy (anonymous):

nope it didnt take it

ganeshie8 (ganeshie8):

take a screenshot and attach

OpenStudy (anonymous):

\[x = \pm \frac{ 1 }{ \sqrt{2} }\] does that work?

OpenStudy (anonymous):

What ganeshie said is correct, maybe rounding errors or something?